OpenStudy (anonymous):

Solve: 4^2x = 32^1/2

OpenStudy (anonymous):

is that (32^1)/2 or 32^1/2?

OpenStudy (anonymous):

it is 32^1/2 (32 exponent 1/2)

OpenStudy (anonymous):

i got x = 1 over 2 square root 2

OpenStudy (anonymous):

Hmm, its not one of my choices. But i did use wolfram alpha nd it got the same.

OpenStudy (anonymous):

what are your choices?

OpenStudy (anonymous):

2/9, 1/2, 3/4, 5/8

OpenStudy (cggurumanjunath):

4^(2x)=(32)^(1/2)

OpenStudy (cggurumanjunath):

(4^((2)^x))=32^(1/2)

OpenStudy (cggurumanjunath):

(4^((2)^x))=32^(1/2) 16^x=32^(1/2)

OpenStudy (cggurumanjunath):

best is take log on both sides !

OpenStudy (anonymous):

So, log16^x = log32^1/2

OpenStudy (phi):

you could write 4 and 2^2 and 32 as 2^5

OpenStudy (cggurumanjunath):

yes

OpenStudy (cggurumanjunath):

x*log 16 =1/2 log(32)

OpenStudy (cggurumanjunath):

x=0.5*(log(32)/log(16))

OpenStudy (phi):

in other words 4^(2x) = 32^½ with 4= 2^2 and 32= 2^5 (2^2)^(2x) = (2^5)^½ which is the same as 2^(4x) = 2^(5/2) now equate exponents (you have the same base)

OpenStudy (cggurumanjunath):

yes ,phi method is also easy.

OpenStudy (anonymous):

then your left with 4x = 5/2 now solve for x by dividing both sides by 4 and get x = 5/8

OpenStudy (cggurumanjunath):

:)

OpenStudy (cggurumanjunath):

yes ! x=5/8

OpenStudy (anonymous):

:) yay!

OpenStudy (cggurumanjunath):

:) (:

OpenStudy (anonymous):

:) thanks @CGGURUMANJUNATH and @phi its starting to make more sense

OpenStudy (phi):

if you are really enthusiastic, try doing the problem CG's way.

OpenStudy (anonymous):

I think that is to hard for me lol :)

OpenStudy (cggurumanjunath):

have u been taught about logarithms ?

OpenStudy (cggurumanjunath):

one can solve this problems by 2 methods ! 1. exponents 2. logarithms

OpenStudy (anonymous):

Yes i have but i dont quite understand it all. some are easy and others are difficult.